A 350 mL solution has an osmotic pressure of 2.73 x 10^-3 atm. If 0.21g of solute was added to the solution, what is the molar mass of the solute?
1
Expert's answer
2020-04-22T09:17:25-0400

Law of VantGoff "\pi=C_m*R*T=((m/M)/V)*R*T"

"M=(m*R*T)/(\pi*V)"

"M=0.21g*R*273,15K*0.35l=(0.21g*0.082*273*0.35)\/mol"

M=602,7g/mol